OpenStudy (anonymous):

Convert the polar equation r^2 = 2sin2Ө to a Cartesian equation. Answer choices: (x^2 + y^2)^2 = 2xy (x^2 + y^2)^2 = 4xy (x^2 + y^2)^2 = 2y^2

OpenStudy (anonymous):

Ok thank you!

OpenStudy (kash_thesmartguy):

Can I get a medal? Just kidding :)

OpenStudy (jtvatsim):

One trick you might want to note for this one is the following fact: \[\sin(2\theta) = 2 \sin(\theta)\cos(\theta)\] That might be useful for you when you try to convert.

OpenStudy (anonymous):

Thank you!

OpenStudy (jtvatsim):

No worries, feel free to repost the question if you get stuck. You should have the basic information you need to succeed now though. :) Good luck!